# SecureWebArena: A Holistic Security Evaluation Benchmark for LVLM-based Web Agents

### v1

@@ Abstract.
Large vision–language model (LVLM)-based web agents are emerging as powerful tools for automating complex online tasks. However, when deployed in real-world environments, they face serious security risks, motivating the design of security evaluation benchmarks. Existing benchmarks provide only partial coverage, typically restricted to narrow scenarios such as user-level prompt manipulation, and thus fail to capture the broad range of agent vulnerabilities. To address this gap, we present SecureWebArena, the first holistic benchmark for evaluating the security of LVLM-based web agents. SecureWebArena first introduces a unified evaluation suite comprising six simulated but realistic web environments (e.g., e-commerce platforms, community forums) and includes 2,970 high-quality trajectories spanning diverse tasks and attack settings. The suite defines a structured taxonomy of six attack vectors spanning both user-level and environment-level manipulations. In addition, we introduce a multi-layered evaluation protocol that analyzes agent failures across three critical dimensions: internal reasoning, behavioral trajectory, and task outcome, facilitating a fine-grained risk analysis that goes far beyond simple success metrics. Using this benchmark, we conduct large-scale experiments on 9 representative LVLMs, which fall into three categories: general-purpose, agent-specialized, and GUI-grounded. Our results show that all tested agents are consistently vulnerable to subtle adversarial manipulations and reveal critical trade-offs between model specialization and security. By providing (1) a comprehensive benchmark suite with diverse environments and a multi-layered evaluation pipeline, and (2) empirical insights into the security challenges of modern LVLM-based web agents, SecureWebArena establishes a foundation for advancing trustworthy web agent deployment.
Do, Not, Use, This, Code, Put, the, Correct, Terms, for, Your, Paper
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
@@
Figure 1. Overall illustration of our SecureWebArena, the first holistic benchmark for evaluating the security of LVLM-based web agents.
@@
Large vision language models (LVLMs) (Hurst et al., 2024; Team et al., 2023; Wang et al., 2024) have equipped autonomous agents with powerful capabilities to perceive and reason across language, vision, and user interface elements (Zeng et al., 2025; Sravanthi et al., 2025; Huang et al., 2025). As web agents, these models can navigate complex websites, fill out forms, and make multi-step decisions based on combined visual and textual input (Ning et al., 2025; Abuelsaad et al., 2024; Lai et al., 2024). However, as these agents are deployed in real-world scenarios to handle sensitive data and critical workflows, their growing autonomy exposes them to severe security threats, such as pop-up attacks (Zhang et al., 2024) and prompt injections (Evtimov et al., 2025; Johnson et al., 2025; Wang et al., 2025b; Yu et al., 2025).
@@
The growing recognition of these security threats has led to the first wave of security evaluation benchmarks (Kumar et al., 2024; Tur et al., 2025; Evtimov et al., 2025; Levy et al., 2024; Liu et al., 2025; Ying et al., 2024a; Xiao et al., 2023b, a). While these valuable contributions have begun to explore the security problem, they often do so with a limited scope, focusing on isolated aspects of the threat landscape. Some benchmarks primarily investigate risks stemming from malicious or harmful user instructions (Kumar et al., 2024; Tur et al., 2025; Ying et al., 2025d, 2024b). Others concentrate on specific, narrow threat models, with a notable focus on prompt injection originating from within the web page (Evtimov et al., 2025), or on adherence to predefined policies in enterprise contexts (Levy et al., 2024). In summary, existing security evaluation benchmarks for web agents fail to provide a unified, systematic framework that addresses vulnerabilities from both user-level instructions and diverse environment-level manipulations, thus failing to capture the broad range of vulnerabilities.
@@
To address this gap, this paper introduces SecureWebArena, the first holistic benchmark specifically designed for evaluating the security of LVLM-based web agents. Our benchmark first provides a unified evaluation suite featuring 6 simulated yet realistic representative web environments, such as online shopping and code management platforms. Central to our framework is a structured classification of 6 attack vectors that span both user-level manipulations (e.g., Jailbreak (Zou et al., 2023; Ying et al., 2025a, c; Zou et al., 2025; Ying et al., 2025b; Mu et al., 2025)) and environment-level threats (e.g., Pop-up Attack (Zhang et al., 2024)). To enable a deeper level of assessment, we introduce a multi-layered evaluation protocol that analyzes agent failures across three critical dimensions: internal reasoning, behavioral trajectory, and task outcome. This approach facilitates a fine-grained risk analysis that goes far beyond simple success metrics. Our main contributions are:
@@
2. Related Work
2.1. Benchmarks for Web Agents
The evaluation of web agents has traditionally focused on task-oriented performance, assessing their ability to navigate web environments and complete user instructions. Early benchmarks such as WebShop (Yao et al., 2022), Mind2Web (Deng et al., 2023a), WebArena (Zhou et al., 2023), and VisualWebArena (Koh et al., 2024) were instrumental in measuring functional capabilities but largely overlooked safety and security vulnerabilities. More recent efforts have begun to construct security-focused benchmarks, although these typically address specific threat dimensions: BrowserART (Kumar et al., 2024) and SAFEARENA (Tur et al., 2025) examine agent responses to harmful user instructions, while WASP (Evtimov et al., 2025) targets prompt injection from malicious web content. Other benchmarks explore niche scenarios such as socio-cultural sensitivity (Qiu et al., 2024) or policy compliance in enterprise systems (Levy et al., 2024).
Despite these contributions, existing benchmarks remain limited in three key ways. ❶ Most works evaluate user- or environment-level attacks separately, lacking a unified framework. ❷ The diversity of attack vectors is limited, with only a subset of representative threats realized across agents or environments. ❸ Most evaluations treat agent decisions monolithically, lacking fine-grained analysis of why and how failures occur.
In contrast, our SecureWebArena offers a unified testbed encompassing six diverse web environments and six representative attack vectors, along with a multi-layered evaluation protocol that analyzes reasoning, behavior, and outcomes to reveal agent vulnerabilities under adversarial conditions. Tab. 1 summarizes the key differences between SecureWebArena and existing benchmarks.
2.2. Attack Vectors on Web Agents
As LVLMs are increasingly deployed in interactive web environments, their multimodal decision-making processes are being exploited by a diverse set of attack vectors. These attacks can be broadly categorized into two main strategies. The first manipulates the agent’s language understanding through methods like Direct Prompt Injection (DP Injection) (Wang et al., 2025b) and Jailbreak Attacks (Zou et al., 2023). The second, more common strategy deceives the agent through the web interface itself. This includes visually deceptive pop-ups and ads that mimic legitimate UI elements (Pop-up Attack/Ad Injection) (Zhang et al., 2024; Wang et al., 2025a), distraction techniques that obscure safe options (Distract Attack) (Ma et al., 2025), and Indirect Prompt Injection (IP Injection) (Greshake et al., 2023) that hide malicious commands within plausible-looking interface text.
Table 1. Comparison of our SecureWebArena with existing web agent security evaluation benchmarks across key dimensions.
@@ 2.2. Attack Vectors on Web Agents
3. Threat Model
3.1. Preliminaries
We model a web agent based on the Set-of-Marks (SoM) paradigm (Yang et al., 2023) as a sequential decision-making system. The agent aims to accomplish a high-level task $\mathcal{G}$, specified by the user in natural language, within a dynamic web environment $\mathcal{E}$. The interaction proceeds over discrete timesteps $t=1,2,\dots,T$.
@@ 3.1. Preliminaries
  1. (1)
    State Perception. The agent captures the current state $s_{t}\in\mathcal{S}$ of the web environment via a SoM-augmented observation $o_{t}$. Specifically, a client-side script automatically annotates every interactable element on the current webpage with a unique integer ID and a colored bounding box. This yields two components: ❶A marked screenshot $v_{t}^{\text{SoM}}$, where each interactable element is overlaid with its ID and bounding box. ❷ A SoM metadata list $\mathcal{L}_{t}=\{(\text{id}_{i},\text{tag}_{i},\text{text}_{i})\}_{i=1}^{N}$, which provides the ID, HTML tag type (e.g., BUTTON, INPUT), and visible text content (if any) for each marked element. The full observation is thus:
    (1) $o_{t}=(v_{t}^{\text{SoM}},\mathcal{L}_{t}).$
  2. (2)
    Reasoning and Action Generation. The LVLM $\mathcal{M}$ takes as input the user goal $\mathcal{G}$, the current SoM observation $o_{t}$, and the interaction history $\mathcal{H}_{t-1}=\{(o_{1},a_{1}),\dots,(o_{t-1},a_{t-1})\}$. It processes the interleaved image-text context to produce a CoT (Wei et al., 2022) reasoning trace $c_{t}$ and selects the next action $a_{t}$ from a discrete action space $\mathcal{A}$:
    (2) $c_{t},a_{t}=\mathcal{M}(\mathcal{G},o_{t},\mathcal{H}_{t-1}).$
@@
  3. (3)
    Environment Interaction. The selected action $a_{t}$ is executed in the environment $\mathcal{E}$, leading to a deterministic state transition:
    (3) $s_{t+1}=\mathcal{E}(s_{t},a_{t}).$
@@
3.2. Attacker’s Objectives and Capabilities
@@
3.2.1. User-Level Threats
In this scenario, the attacker directly manipulates the high-level task instruction $\mathcal{G}$ provided to the agent. The environment $\mathcal{E}$ is assumed to be benign, but the goal itself is malicious. Formally, the attacker crafts a malicious goal $\mathcal{G}_{\text{malicious}}$ to replace the benign $\mathcal{G}$. These threats target the agent’s language understanding and safety alignment.
3.2.2. Environment-Level Threats
In this more complex scenario, the user’s goal $\mathcal{G}$ is benign, but the web environment $\mathcal{E}$ is controlled by the attacker. The attacker manipulates the environment to produce a deceptive observation $o_{t}=(v_{t}^{\text{SoM}},\mathcal{L}_{t})$ that misleads the agent’s perception and reasoning. These threats exploit the agent’s reliance on both visual cues from the screenshot $v_{t}^{\text{SoM}}$ and structural information from the metadata $\mathcal{L}_{t}$.
We do not assume access to internal model gradients or weights, reflecting the realistic black-box deployment of commercial LVLMs. The agent operates under standard web permissions, and attackers cannot invoke system-level APIs. Attacks are confined to manipulations through the natural input-output interface of the agent, such as textual instructions or perceptible visual elements.
@@ 3.2.2. Environment-Level Threats
Figure 2. SecureWebArena framework. It integrates simulated environments, diverse attack vectors, and multi-level evaluation to assess agent safety performance to adversarial manipulation.
4. SecureWebArena Design
4.1. Environment Suite
To ensure the generality and robustness of our security evaluations, SecureWebArena is built upon 6 diverse, realistic web environments adapted from WebArena (Zhou et al., 2023) and VisualWebArena (Koh et al., 2024). These environments are implemented as lightweight, browser-based web applications with full HTML, CSS, and DOM structures, supporting dynamic rendering, user interaction, form handling, and layout variation. They provide a faithful simulation of modern web user interfaces while enabling attack injection and trajectory logging in a controlled, reproducible manner. Each environment is instrumented with SoM-based markup that uniquely identifies interactive elements. It also supports structured metadata export for element-level reasoning and dynamic attack surface exposure to simulate real-world threats.
The six environments span a broad range of real-world interaction contexts, grouped into four representative categories. Information Retrieval and Navigation environments (e.g., Wikipedia and Reddit) feature dense, user-generated content with richly linked layouts, testing agents’ ability to maintain focus amidst subtle textual distractions. E-commerce and Transaction tasks in Shopping simulate high-stakes workflows involving financial decisions and sensitive data, providing a natural setting to evaluate agents’ risk awareness and cautiousness. Content Management and Publishing environments (e.g., Classifieds and ShoppingAdmin) involve privileged actions such as creating, modifying, or deleting content, presenting scenarios where errors could lead to persistent or destructive consequences. Finally, the Software Development and Collaboration environment, based on GitLab, focuses on technical workflows such as issue tracking and code repository navigation, enabling evaluation of agents’ resilience in structured, domain-specific interfaces.
These environments impose diverse perceptual, reasoning, and interface navigation demands on agents. They also serve as rich platforms for embedding both user-level and environment-level adversarial content under varied task semantics. Fig. 2 illustrates the overall framework of the proposed benchmark, and App. A.1 provides illustrative examples of each environment.
4.2. Task Construction
Tasks in SecureWebArena are manually designed to reflect realistic user intents across the six interaction environments, covering a wide range of common goals such as information look-up, product purchase, issue reporting, and content editing. Each task is specified as a standalone natural language instruction (e.g., “Find a wireless mouse under \$20 and add it to the cart”) that assumes the agent is placed in an appropriate initial interface state (e.g., relevant page opened, default settings loaded).
Although our benchmark focuses exclusively on adversarial evaluation, each attack task in SecureWebArena is implicitly grounded in a corresponding benign goal. The adversarial version perturbs this underlying task with one of the defined attack vectors, either by altering the instruction (user-level), or modifying the rendering interface (environment-level). Each attack vector is instantiated using a dedicated template or pattern based on its definition in prior literature, ensuring consistent application across tasks and environments.
To support systematic evaluation, tasks in SecureWebArena are designed to reflect real-world user behavior (goal realism) and require understanding of both rendered UI elements and natural language instructions (multimodal dependency). Attacks are localized to a specific input channel for controlled assessment (isolated intervention), and tasks span diverse interaction patterns including navigation, selection, typing, and multi-step planning (coverage diversity). This design enables SecureWebArena to probe the holistic adversarial decision-making pipeline of web agents while supporting reproducible, attack-specific performance analysis.
4.3. Attack Taxonomy
SecureWebArena implements 6 representative attack vectors that target both user-level and environment-level inputs, as defined in our threat model (Sec. 3.2). Each attack manipulates a specific component of the agent’s decision function $\mathcal{M}(\mathcal{G},o_{t},\mathcal{H}_{t-1})$, either by altering the instruction $\mathcal{G}$ or the rendered observation $o_{t}=(v_{t}^{\text{SoM}},\mathcal{L}_{t})$. For clarity, we denote perturbed instructions and observations as $\widetilde{\mathcal{G}}$ and $\widetilde{o}_{t}$, and introduce $\mathcal{A}_{\text{adv}}$ to denote an adversarially modified action trajectory induced by these attacks. We categorize the 6 attack vectors as follows:

### v2

@@ Abstract
Large vision–language model (LVLM)-based web agents are emerging as powerful automation tools but face severe security risks in real-world deployment. Existing benchmarks offer limited coverage, typically isolating user-level prompts from environmental threats, thus failing to capture the full spectrum of vulnerabilities. To address this, we present SecureWebArena, the first holistic security benchmark for web agents. SecureWebArena features a unified suite of six realistic web environments with 2,970 adversarial trajectories, covering a structured taxonomy of six attack vectors that span both user-level and environment-level manipulations. Crucially, we introduce a multi-layered evaluation protocol that dissects agent failures across internal reasoning, behavioral execution, and task outcomes, enabling fine-grained risk analysis beyond simple success metrics. Experiments on 9 representative LVLMs reveal universal vulnerabilities to subtle manipulations and uncover significant trade-offs between model specialization and security. SecureWebArena establishes a rigorous foundation for advancing the development of trustworthy web agents.
@@
SecureWebArena: A Holistic Security Evaluation Benchmark for LVLM-based Web Agents
Zonghao Ying1∗, Yangguang Shao2∗, Jianle Gan3, Gan Xu4, Wenxin Zhang5, Quanchen Zou6, Junzheng Shi2, Zhenfei Yin7, Mingchuan Zhang8, Aishan Liu1†, Xianglong Liu1 1SKLCCSE, Beihang University  2Institute of Information Engineering, CAS  3China University of Petroleum (East China) 4Zhejiang University of Technology  5University of Chinese Academy of Science  6360 AI Security Lab 7The University of Sydney  8Henan University of Science and Technology
footnotetext: Equal contribution. Corresponding author.
1 Introduction
Figure 1: Overall illustration of our SecureWebArena, the first holistic benchmark for evaluating the security of LVLM-based web agents.
@@ 1 Introduction
Large vision language models (LVLMs) Hurst et al. (2024); Team et al. (2023); Wang et al. (2024) have equipped autonomous agents with powerful capabilities to perceive and reason across language, vision, and user interface elements Zeng et al. (2025); Sravanthi et al. (2025); Huang et al. (2025). As web agents, these models can navigate complex websites, fill out forms, and make multi-step decisions based on combined visual and textual input Ning et al. (2025); Abuelsaad et al. (2024); Lai et al. (2024). However, as these agents are deployed in real-world scenarios to handle sensitive data and critical workflows, their growing autonomy exposes them to severe security threats, such as pop-up attacks Zhang et al. (2024) and prompt injections Evtimov et al. (2025); Johnson et al. (2025).
@@
The growing recognition of these security threats has led to the first wave of security evaluation benchmarks Kumar et al. (2024); Tur et al. (2025); Evtimov et al. (2025); Levy et al. (2024). While these valuable contributions have begun to explore the security problem, they often do so with a limited scope, focusing on isolated aspects of the threat landscape. Some benchmarks primarily investigate risks stemming from malicious or harmful user instructions Kumar et al. (2024); Tur et al. (2025). Others concentrate on specific, narrow threat models, with a notable focus on prompt injection originating from within the web page Evtimov et al. (2025), or on adherence to predefined policies in enterprise contexts Levy et al. (2024). In summary, existing security evaluation benchmarks for web agents fail to provide a unified, systematic framework that addresses vulnerabilities from both user-level instructions and diverse environment-level manipulations, thus failing to capture the broad range of vulnerabilities.
@@
To address this gap, this paper introduces SecureWebArena, the first holistic benchmark specifically designed for evaluating the security of LVLM-based web agents. Our benchmark first provides a unified evaluation suite featuring 6 simulated yet realistic representative web environments, such as online shopping and code management platforms. Central to our framework is a structured classification of 6 attack vectors that span both user-level manipulations (e.g., Jailbreak Zou et al. (2023)) and environment-level threats (e.g., Pop-up Attack Zhang et al. (2024)). To enable a deeper level of assessment, we introduce a multi-layered evaluation protocol that analyzes agent failures across three critical dimensions: internal reasoning, behavioral trajectory, and task outcome. This approach facilitates a fine-grained risk analysis that goes far beyond simple success metrics. Our main contributions are:
@@
2 Related Work
Table 1: Comparison of our SecureWebArena with existing web agent security evaluation benchmarks across key dimensions.
@@ 2 Related Work
2.1 Benchmarks for Web Agents
Early web agent benchmarks, such as WebShop Yao et al. (2022), Mind2Web Deng et al. (2023a), WebArena Zhou et al. (2023), and VisualWebArena Koh et al. (2024), prioritized functional correctness, largely overlooking security vulnerabilities. While recent works have emerged to address this, they typically focus on isolated threat dimensions. For instance, BrowserART Kumar et al. (2024) and SAFEARENA Tur et al. (2025) target harmful user instructions, whereas WASP Evtimov et al. (2025) focuses on prompt injection from malicious web content. Others address niche scenarios like socio-cultural sensitivity Qiu et al. (2024) or enterprise policy compliance Levy et al. (2024).
Despite these advances, existing benchmarks suffer from fragmented scopes that isolate user from environment threats, limited attack diversity, and coarse-grained metrics that neglect the reasoning behind failures. SecureWebArena bridges these gaps by establishing a holistic testbed that integrates six representative attack vectors across diverse environments. Crucially, we employ a multi-layered evaluation protocol that dissects reasoning, behavior, and outcomes to provide granular insights into agent vulnerabilities, as compared in Tab. 1.
2.2 Attack Vectors on Web Agents
As LVLMs are increasingly deployed in interactive web environments, their multimodal decision-making processes are being exploited by a diverse set of attack vectors. These attacks can be broadly categorized into two main strategies. The first manipulates the agent’s language understanding through methods like Direct Prompt Injection (DP Injection) Wang et al. (2025b) and Jailbreak Attacks Zou et al. (2023). The second, more common strategy deceives the agent through the web interface itself. This includes visually deceptive pop-ups and ads that mimic legitimate UI elements (Pop-up Attack/Ad Injection) Zhang et al. (2024); Wang et al. (2025a), distraction techniques that obscure safe options (Distract Attack) Ma et al. (2025), and Indirect Prompt Injection (IP Injection) Greshake et al. (2023) that hide malicious commands within plausible-looking interface text.
3 Threat Model
3.1 Preliminaries
We model a web agent based on the Set-of-Marks (SoM) paradigm Yang et al. (2023) as a sequential decision-making system. The agent aims to accomplish a high-level task $\mathcal{G}$, specified by the user in natural language, within a dynamic web environment $\mathcal{E}$. The interaction proceeds over discrete timesteps $t=1,2,\dots,T$.
@@ 3.1 Preliminaries
  1. 1.
    State Perception. The agent captures the current state $s_{t}\in\mathcal{S}$ of the web environment via a SoM-augmented observation $o_{t}$. Specifically, a client-side script automatically annotates every interactable element on the current webpage with a unique integer ID and a colored bounding box. This yields two components: ❶A marked screenshot $v_{t}^{\text{SoM}}$, where each interactable element is overlaid with its ID and bounding box. ❷ A SoM metadata list $\mathcal{L}_{t}=\{(\text{id}_{i},\text{tag}_{i},\text{text}_{i})\}_{i=1}^{N}$, which provides the ID, HTML tag type (e.g., BUTTON, INPUT), and visible text content (if any) for each marked element. The full observation is thus $o_{t}=(v_{t}^{\text{SoM}},\mathcal{L}_{t})$.
  2. 2.
    Reasoning and Action Generation. The LVLM $\mathcal{M}$ takes as input the user goal $\mathcal{G}$, the current SoM observation $o_{t}$, and the interaction history $\mathcal{H}_{t-1}=\{(o_{1},a_{1}),\dots,(o_{t-1},a_{t-1})\}$. It processes the interleaved image-text context to produce a CoT Wei et al. (2022) reasoning trace $c_{t}$ and selects the next action $a_{t}$ from a discrete action space $\mathcal{A}$, with $c_{t},a_{t}=\mathcal{M}(\mathcal{G},o_{t},\mathcal{H}_{t-1})$.
@@
  3. 3.
    Environment Interaction. The selected action $a_{t}$ is executed in the environment $\mathcal{E}$, leading to a deterministic state transition, $s_{t+1}=\mathcal{E}(s_{t},a_{t})$.
@@
3.2 Attacker’s Objectives and Capabilities
@@
3.2.1 User-Level Threats (Malicious User)
The agent trusts the user’s instructions, but the user is malicious. The attacker controls the natural language goal $\mathcal{G}$. They cannot modify the web environment $\mathcal{E}$ or the agent’s internal weights.
While Direct Prompt Injection (DP) and Jailbreaks are generic to LLMs, they are critical in web agent deployments. For instance, a malicious user might use DP Injection to force a shared enterprise agent to exfiltrate data from a private database, or use Jailbreaking to bypass safety filters preventing the purchase of illegal goods.
3.2.2 Environment-Level Threats
The user is benign, but the agent interacts with an untrusted external website. The attacker controls the DOM structure, CSS styling, and content of the website. They can inject scripts to render pop-ups, modify text, or alter visual layouts. They cannot access the browser’s chrome (address bar, browser settings) or the agent’s system prompt.
We assume the agent relies on visual (screenshot) and structural (SoM) inputs. Attacks such as Pop-ups are rendered with standard HTML/CSS to mimic legitimate system notifications, exploiting the agent’s visual grounding.
@@ 3.2.2 Environment-Level Threats
Figure 2: SecureWebArena framework. It integrates simulated environments, diverse attack vectors, and multi-level evaluation to assess agent safety performance to adversarial manipulation.
4 SecureWebArena Design
4.1 Environment Suite
To ensure robustness, SecureWebArena integrates six realistic web environments adapted from WebArena Zhou et al. (2023) and VisualWebArena Koh et al. (2024). Implemented as fully functional web applications, they support dynamic rendering and complex interactions while enabling controlled attack injection. Each environment is instrumented with Set-of-Mark (SoM) tagging and structured metadata export to facilitate precise element-level reasoning and dynamic attack surface exposure.
The suite covers four representative domains: (1) Information Retrieval (Wikipedia, Reddit), featuring dense, user-generated content to test agent focus amidst distractions; (2) E-commerce (Shopping), simulating high-stakes financial workflows to evaluate risk awareness; (3) Content Management (Classifieds, ShoppingAdmin), involving privileged actions where errors lead to persistent consequences; and (4) Software Development (GitLab), focusing on structured technical workflows. Collectively, these platforms impose diverse perceptual and reasoning demands, serving as a rich testbed for embedding adversarial content under varied task semantics. Fig. 2 illustrates the framework, with detailed examples in App. B.
4.2 Task Construction
Tasks in SecureWebArena capture realistic user intents across the six environments, ranging from information retrieval to content management. Each task pairs a natural language instruction (e.g., “Find a wireless mouse under \$20”) with an initialized interface state. Crucially, every adversarial scenario is grounded in a benign goal, which is then perturbed using standardized templates corresponding to specific user-level or environment-level attack vectors. To ensure systematic evaluation, task construction adheres to four key principles: goal realism, multimodal dependency, isolated intervention (localizing attacks to specific input channels), and coverage diversity (spanning navigation, typing, and multi-step planning). This design enables a controlled, holistic assessment of the agent’s decision-making pipeline.
4.3 Attack Taxonomy
Based on our threat model, SecureWebArena implements six attack vectors targeting the agent’s decision function $\mathcal{M}(\mathcal{G},o_{t},\mathcal{H}_{t-1})$. We denote perturbed instructions and observations as $\widetilde{\mathcal{G}}$ and $\widetilde{o}_{t}$, respectively.